OpenStudy (anonymous):

is it possible to solve 0 / 51

OpenStudy (australopithecus):

yes

OpenStudy (anonymous):

0/anything = 0

OpenStudy (australopithecus):

except 0/0 which is undefined

OpenStudy (anonymous):

No. 0/Any number = 0 And, Any number/0 = Infinite.

OpenStudy (australopithecus):

@Fall12-13 You cannot divide by 0, it is undefined

OpenStudy (anonymous):

okay. i confuse between the 0 because one of them is impossible to solve. Thanks.

OpenStudy (australopithecus):

Just to be clear any number/0 =/= infinite it is undefined

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got it now.

OpenStudy (anonymous):

@Australopithecus http://www.youtube.com/watch?v=DkrqVgb-his

OpenStudy (anonymous):

i always tend to think things divided by 0 should be 0. If i divide something into 2 pieces i have 2, if i divide it into 1 piece i have one. if i divide something into no pieces i have no pieces i have 0

OpenStudy (anonymous):

You can also think this way. Anything divided by 0 can be divided infinite number of times. For example: If I have 4 candies, and I divide them to 2 kids, I can only at the maximum give 2 to each one of them. But suppose I have 4 candies and there are no kids at all, then I can divide them indefinitely because there is no real division going on.

OpenStudy (anonymous):

a good point. what about \[4 \div \frac{1}{0}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

i think

OpenStudy (anonymous):

i thought it would be \[\frac{4 \times 0}{1} = 0\]

OpenStudy (anonymous):

It would be infinity. Any number divided by infinity is infinity because then the denominator changes to 0.

OpenStudy (anonymous):

ok